Two dice are tossed 288 times. How many times would you expect to get a sum of 4?
Answer by mananth(16949)

when two dice are rolled there are 36 possibilities
getting a sum of 4 (1,3),(2,2) (3,1)
p(A) = 3/36 =1/12
it is rolled 288 times
Number of times you can expect = 1/12 * 288 = 24